OpenStudy (anonymous):

Solve the given system of equations using addition or substraction. x+y=8 & x-y=2

OpenStudy (anonymous):

\[x + y = 8\]\[x - y = 2\]Solve by elimination. \[2x = 10\]\[x = 5\]Now plug it back in and solve for y. \[x + y = 8\]\[5 + y = 8\]Solve for y and you're done :)

OpenStudy (anonymous):

Thanks a million!

OpenStudy (anonymous):

np :)